# Asynchronous Verified Semantic Caching for Tiered LLM Architectures

### v1

@@ Abstract
[…]
[Correspondence]Asmit Kumar Singh: asingh77@apple.com; Haozhe Wang: haozhe_wang@apple.com; Santosh Attaluri: lattaluri@apple.com; Tak Chiam: t_chiam@apple.com;Weihua Zhu: weihua_zhu@apple.com
1 Introduction
@@
(b) Krites asynchronous verified semantic caching policy.
Figure 1: (a) Hit rate composition for the baseline static threshold policy versus Krites. The overall cache hit rate (total bar height) remains identica. However, Krites significantly increases the proportion of hits served directly from the offline vetted, curated static tier (darker bottom section), strictly improving the safety and quality of served responses.(b) Krites couples the static and dynamic tiers via an off the path LLM judge and auxiliary overwrite, building on the static–dynamic architectures used in web search (Fagni et al., 2006; Baeza-Yates et al., 2008; Mele et al., 2020).
@@
Incoming query $q$ Cached neighbor $h$ Similarity $s$
What’s the word on my dog having honey Can my dog have honey 0.87
Was anyone successful in the lottery last night anybody win the lottery last night 0.93
@@
2 Background and system model
@@
2.1 Vector embeddings and semantic search
@@
$\operatorname{sim}(q,x)=\frac{v_{q}\cdot v_{x}}{\|v_{q}\|\,\|v_{x}\|}.$ (2.2)
@@
2.2 Tiered static-dynamic architecture
@@
2.2.1 Static tier $C_{\text{static}}$
@@
2.2.2 Dynamic tier $C_{\text{dynamic}}$
@@ 2.2.3 Agentic backend $B$
2.3 Baseline GPTCache style policy
@@ 3 Krites architecture
15:  enqueue asynchronous task $\textsc{VerifyAndPromote}(q,h_{\text{static}},v_{q})$
@@
3.1 Grey-zone trigger and task scheduling
@@ 3.2 LLM judge requirements
The function J(q,h,a) requires a binary semantic judge: given an incoming prompt q, a cached prompt h, and the cached answer a, it must decide whether a is an acceptable answer for q. In a production setting, this can be implemented via a strict, rubric guided LLM prompt evaluating intent, entity matching, and freshness constraints. Because such LLM-as-a-judge formulations already achieve high agreement with semantic equivalence labels (Zheng et al., 2023; Gu et al., 2024; Tan et al., 2024), our evaluation (Section 4) abstracts the specific prompt implementation and instead uses an oracle derived from ground truth equivalence classes to establish the policy’s model agnostic performance bounds.
3.3 Auxiliary overwrite semantics
@@
3.4 Cost and capacity considerations
@@
4 Experimental evaluation
We evaluate Krites with trace-driven simulations on the two open benchmarks introduced by vCache (Schroeder et al., 2025). In all experiments we reuse the embeddings, ground truth equivalence classes. Krites requires a binary decision $J(q,h,a)$ to approve promotions. To keep our evaluation model-agnostic and consistent with vCache, we do not run an LLM judge during simulation. Instead, we instantiate $J$ directly from the benchmark equivalence relation: we approve iff the query $q$ and the candidate neighbor $h$ share the same ground truth class under the benchmark labels. Notably, the dataset labels were originally generated through LLM judges.
4.1 Datasets and static-dynamic split
@@ 4.3 Results
Table 1: Static-origin served fraction for the tuned static baseline and Krites. For the baseline this equals the direct static tier hit rate; for Krites it includes both direct static tier hits and dynamic tier hits created by auxiliary overwrite promotions. The baseline thresholds are taken from the Pareto optimal GPTCache configurations reported in Schroeder et al. (2025).
(a) SemCacheLMArena Static-Cache Hit Rate.
(b) SemCacheSearchQueries Static-Cache Hit Rate.
Figure 2: Static origin served fraction (including auxiliary-overwrite promotions) as a function of requests processed, starting from a cold dynamic cache, for (a) SemCacheLMArena and (b) SemCacheSearchQueries. Krites increases static-origin coverage over time by populating the dynamic tier with verified pointers to static answers via auxiliary overwrites.
@@ 5 Discussion
Our trace-driven evaluation instantiates $J$ as an oracle defined by the vCache benchmark equivalence labels. The reported gains therefore quantify the benefit of asynchronous verification assuming a high precision verifier that matches the benchmark’s interchangeability notion. In production, an LLM-based verifier may have nonzero false reject and false approve rates. False rejects reduce promotion volume (reducing gains), while false approves can introduce additional errors on promoted entries. Since verification is off path, such errors do not affect the request that triggered verification, but can affect subsequent hits to promoted keys; if promoted hit traffic is $p_{\text{prom}}$ and the verifier’s false-approve rate on judged pairs is $\epsilon$, then the incremental error contribution from promotions is upper-bounded by $\epsilon\,p_{\text{prom}}$.
5.1 Cost and ROI of off-path judging
Krites introduces a new resource tradeoff: background judge compute in exchange for higher static origin coverage. Let $\lambda$ be the request rate and let $p_{\text{grey}}$ be the fraction of requests whose nearest static neighbor falls in the grey zone $\sigma_{\min},\tau_{\text{static}})$. The judge request rate is approximately $\lambda_{J}\approx\lambda\,p_{\text{grey}}$ (before deduplication). Let $c_{J}$ denote the average cost (or latency/compute) per judge call.
The benefit of a single judge approval depends on (i) whether the pair is approved and (ii) how many times the promoted key is reused before eviction. Let $p_{\text{app}}$ be the approval probability and let $N$ be the number of subsequent requests that hit the promoted dynamic entry before eviction. Then the expected number of additional static origin serves attributable to one judge call is $\mathbb{E}[p_{\text{app}}N]$. This quantity is workload-dependent: it is larger when paraphrases recur and when the dynamic cache retains promoted entries long enough for reuse.
In practice, systems can improve ROI by (i) deduplicating repeated $(q,h_{\text{static}})$ judgments, (ii) gating judgment on observed recurrence (e.g., only judge when $q$ has appeared multiple times in a short window), and (iii) rate limiting the judge pool to a fixed budget while prioritizing higher frequency or higher value prompts. These controls make Krites compatible with strict compute budgets while preserving its iso-latency serving behavior.
Apple and the Apple logo are trademarks of Apple Inc., registered in the U.S. and other countries and regions.

### v2

@@ Abstract.
semantic caching, LLM serving systems, tiered caching, LLM-as-a-judge, systems
copyright: nonejournalyear: 2026
1. Introduction
@@
(b) Krites policy.
Figure 1. (a) Hit rate composition for the baseline static threshold policy versus Krites. The overall cache hit rate (total bar height) remains identical. However, Krites increases the proportion of requests served with curated static-origin answers (direct static hits plus verified promotions). (b) Krites couples the static and dynamic tiers via an off-path LLM judge and auxiliary overwrite.
Left: stacked bars comparing baseline vs. Krites hit composition. Right: two-tier cache diagram with off-path LLM judge and auxiliary overwrite into the dynamic cache.
@@
Incoming query $q$ Cached neighbor $h$
What’s the word on my dog having honey? Can my dog have honey?
Was anyone successful in the lottery last night? Anybody win the lottery last night?
@@
2. Background and system model
@@
2.1. Vector embeddings and semantic search
@@
(2) $\operatorname{sim}(q,x)=\frac{v_{q}\cdot v_{x}}{\|v_{q}\|\,\|v_{x}\|}.$
@@
2.2. Tiered static-dynamic architecture
@@
2.2.1. Static tier $C_{\text{static}}$
@@
2.2.2. Dynamic tier $C_{\text{dynamic}}$
@@ 2.2.3. Agentic backend $B$
2.3. Baseline GPTCache style policy
@@ 3. Krites architecture
[…]
@@
3.1. Grey-zone trigger and task scheduling
@@ 3.2. LLM judge requirements
The function $J(q,h,a)$ is a binary compatibility judge: given an incoming prompt $q$, a cached prompt $h$, and the cached answer $a$, it decides whether serving $a$ for $q$ is acceptable. In deployment, $J$ can be implemented with a strict, rubric-guided LLM prompt that checks intent alignment, entity/constraint consistency (names, locations, time, quantities), and freshness/personalization requirements, and returns a single-token decision (APPROVE/REJECT) at temperature 0.
Oracle for judge-agnostic policy evaluation.
Our main trace-driven results (Section 4) are intentionally judge-agnostic: we instantiate $J$ as an oracle derived from the benchmark equivalence classes. This isolates the caching policy from judge prompt/model choices and avoids conflating judge behavior with occasional inconsistencies we observed in the benchmark IDs when manually auditing near-duplicate prompts.
Sanity checking judge realism.
To ground practicality, we additionally evaluated a concrete LLM judge (Claude Opus 4.5) on a 100-pair human-audited sample drawn from the static similarity grey zone (i.e., pairs with $s_{\text{static}}\in[\sigma_{\min},\tau_{\text{static}})$). Using the same strict rubric as above and a single-token output format, the judge agreed with human accept/reject labels on 99/100 pairs. This supports that high-precision LLM-based verification is feasible in practice, while Krites keeps verification off the critical path.
3.3. Auxiliary overwrite semantics
@@
3.4. Cost and capacity considerations
@@
4. Experimental evaluation
We evaluate Krites with trace-driven simulations on the two open benchmarks introduced by vCache (Schroeder et al., 2025). In all experiments we reuse the embeddings and the benchmark equivalence classes. Krites requires a binary decision $J(q,h,a)$ to approve promotions. To keep the evaluation policy-centric, judge-agnostic, and consistent with vCache, we do not run an LLM judge inside the simulation loop. Instead, we instantiate $J$ directly from the benchmark equivalence relation: we approve iff the query $q$ and the candidate neighbor $h$ share the same labeled equivalence class. To show the efficacy of current LLMs as judge we follow 3.2
4.1. Datasets and static-dynamic split
@@
Why an oracle is a realistic proxy.
The vCache benchmarks define interchangeability using LLM-as-a-judge labeling procedures, and SemCacheSearchQueries equivalence classes are constructed via a union-find pass guided by an LLM judge (GPT-4.1-nano). Using the benchmark equivalence relation as $J$ therefore evaluates Krites under the benchmark’s own definition of correctness and provides an upper bound for any production verifier. However, in a small manual audit we observed occasional inconsistencies in the benchmark IDs for near-duplicate prompts, which motivates keeping our headline policy results judge agnostic rather than tying them to any single verifier implementation. To sanity check verifier realism, we additionally evaluated Claude Opus 4.5 as a strict binary judge on a 100 pair human-audited sample drawn from the static similarity grey zone, and found 99/100 agreement with human accept/reject labels.
@@ 4.3. Results
Table 1. Static-origin served fraction for the tuned static baseline and Krites. For the baseline this equals the direct static tier hit rate; for Krites it includes both direct static tier hits and dynamic tier hits created by auxiliary overwrite promotions. The baseline thresholds are taken from the Pareto optimal GPTCache configurations reported in Schroeder et al. (2025).
(a) SemCacheLMArena static-origin served fraction over time.
(b) SemCacheSearchQueries static-origin served fraction over time.
Figure 2. Static-origin served fraction (direct static hits plus verified promotions) as a function of requests processed, starting from a cold dynamic cache, for (a) SemCacheLMArena and (b) SemCacheSearchQueries. Krites increases static-origin coverage over time by populating the dynamic tier with verified pointers to static answers via auxiliary overwrites.
Two line plots showing the cumulative static-origin served fraction over processed requests for SemCacheLMArena and SemCacheSearchQueries, comparing Krites to a static threshold baseline.
@@ 5. Discussion
Our trace-driven evaluation instantiates $J$ as an oracle defined by the vCache benchmark equivalence labels, so the headline gains characterize the policy under an ideal verifier for the benchmark notion of interchangeability. To assess realism, we additionally evaluated Claude Opus 4.5 as a strict binary judge on a 100-pair human-audited sample from the similarity grey zone and found 99/100 agreement with human accept/reject labels. In production, an LLM-based verifier may still have nonzero false reject and false approve rates. False rejects reduce promotion volume (reducing gains), while false approves can introduce errors on promoted entries. Since verification is off path, such errors do not affect the request that triggered verification, but can affect subsequent hits to promoted keys; if promoted hit traffic is $p_{\text{prom}}$ and the verifier’s false-approve rate on judged pairs is $\epsilon$, then the incremental error contribution from promotions is upper-bounded by $\epsilon\,p_{\text{prom}}$.
5.1. Cost and ROI of off-path judging
Krites introduces a new resource tradeoff: background judge compute in exchange for higher static-origin coverage. Let $\lambda$ be the request rate and let $p_{\text{grey}}$ be the fraction of requests whose nearest static neighbor falls in the grey zone $\sigma_{\min},\tau_{\text{static}})$. The raw judge request rate is approximately $\lambda_{J}\approx\lambda\,p_{\text{grey}}$ (before deduplication/rate limiting). Let $c_{J}$ denote the average compute cost per judge call.
The benefit of a single judge approval depends on (i) whether the pair is approved and (ii) how many times the promoted key is reused before eviction. Let $p_{\text{app}}$ be the approval probability and let $N$ be the number of subsequent requests that hit the promoted dynamic entry before eviction. Then the expected number of additional static-origin serves attributable to one judge call is $\mathbb{E}[p_{\text{app}}N]$. This quantity is workload-dependent: it is larger when paraphrases recur and when the dynamic cache retains promoted entries long enough for reuse.
Model choice.
Using a large model (e.g., Claude Opus 4.6) as the judge can be compute expensive because the judge prompt includes the cached answer, but it does not affect critical path latency. If compute budget is tight, operators can use a smaller judge model and/or gate judging on recurrence (e.g., only judge after observing a repeated query), trading off promotion volume and verifier precision/recall. In practice, the best verifier choice is application and workload dependent; the oracle-based evaluation in Section 4 isolates Krites from these model-specific details.